# Learning Vision-Based Neural Network Controllers with Semi-Probabilistic Safety Guarantees

### v1

@@ V-C Adaptive Training Data
After we have enough datapoints in $S_{A}$, future training batch $T$ with size $L$ consists of $\lfloor p\cdot L\rfloor$ datapoints from $S_{A}$ and the rest from $S_{0}$, where $p$ is a tunable parameter during training. The portion of datapoints from $S_{0}$ can be sampled uniformly. The portion of datapoints from $S_{A}$ are sampled following the Efraimidis & Spirakis [36] weighted sampling approach to prioritize datapoints that are more difficult (less safe) than others. Specifically, for each datapoint $i$ in $S_{A}$, we first calculate a safety parameter
$z^{(i)}$ $=\mathcal{L}_{safety}^{(i)}$ (6)
$+\gamma\Big{(}\max(0,|\overline{\sigma}^{(i)}_{K}|-\beta)+\max(0,|\underline{\sigma}^{(i)}_{K}|-\beta)\Big{)},$
where the second term further penalizes datapoints whose upper and lower bound is outside of the $[-\beta,\beta]$ region, where $\beta$ is a predefined safety threshold, and ensure such difficult-to-verify inputs are more likely to be sampled. We then assign a weight $w_{i}=e^{\alpha\cdot z^{(i)}}$ (where $\alpha$ is a hyperparameter) to each datapoint $i$ and normalize the weights such that $\sum_{i=1}^{N}w_{i}=1$. Finally we assign each element $x_{i}$ a key $k_{i}=w_{i}^{1/U_{i}}$, where $U_{i}\sim\text{Uniform}(0,1)$. Selecting the top $L$ elements with the highest $k_{i}$ values yields a weighted random sample without replacement, where each element $x_{i}$ is included with probability: $P(x_{i}\in S)=w_{i}/\sum_{j\in\mathcal{B}}w_{j}$.
The full algorithm for data sampling is shown in Algorithm 1. Line $3$ calculates how many datapoints to sample from the buffer based on a tunable parameter $p$; line $4-8$ computes the weights and keys following [36]; line $9-11$ construct the training batch. This approach enables us to have an adaptive training set which maintains the controller’s average performances while improving its verifiability with respect to datapoints that are more difficult.
@@
5:     Compute $z^{(i)}$ according to Equation 6
6:     Compute weight $w_{i}=e^{\alpha\cdot z^{(i)}}$
@@
9:  $T_{A}\leftarrow$ top $L_{A}$ elements in $S_{A}$ by $k_{i}$ values
10:  $T_{0}\xleftarrow{}$ uniformly sample $L_{0}$ points from $\mathcal{S}_{0}$
11:  $T\leftarrow T_{A}\cup T_{0}$
@@ V-D Curriculum Learning
Finally, as in prior work [8], we use curriculum learning. Specifically, during training, we first target $K_{i}$-step verified safety. After a series of epochs, we progress to $K_{i+1}$-step verified safety, where $1\leq K_{1}<K_{2}<\cdots<K_{n-1}<K_{n}=K$. This gradual increase in the verification horizon helps improve training stability and enables the controller to learn more complex safety constraints.
VI Experiments
VI-A Experiment Setup
@@
We evaluate our approach in three settings: 1) the X-Plane 11 Flight Simulator, 2) the CARLA Simulator, and 3) a mini-city miniature urban physical autonomous driving platform with an F1Tenth racing car. We use two evaluation metrics: 1) empirical performance and 2) lower bound probability for safety guarantee as a function of $K$. We use three baselines for comparison: 1) RESPO [27], a safe reinforcement learning framework using iterative reachability analysis; 2) SAC-RCBF [9], which incorporates safety as a robust-control-barrier-function layer into training; and 3) VSRL [8], which guarantees finite-horizon safety by integrating incremental reachability verification into safe reinforcement learning.
@@
Specifically, we first consider the autonomous aircraft taxiing problem using the X-Plane 11 Flight Simulator [37]. For training the image generator, we collected 20,000 state-image pairs. Each sample consists of an RGB image captured by a forward-facing camera, the corresponding state information (lateral offset $d$ ranging from $-10$ to $10$ meters, heading error $\theta$ between $-0.5$ and $0.5$ radians), and control inputs (nose wheel steering angle $\gamma$). Data was collected while the aircraft operated at a constant ground speed.
@@
Our second set of experiments consider autonomous lane following using the CARLA Simulator [38] version 0.9.14. For training the image generator, we sample initial states consisting of the lateral distance $d$ from the lane center, the heading error $\theta$ relative to the lane direction, and the global coordinates $(x,y,z)$ within the CARLA map. The dataset includes trajectories with $d\in[-0.8,0.8]$ meters and $\theta\in[-0.15,0.15]$ radians. Final dataset contains 20,000 state-image pairs collected across different towns (maps) and environmental conditions.
@@
Finally, we evaluated our approach using the F1Tenth racing car, an open-source 1:10 scale autonomous vehicle, for lane following in the mini-city physical testbed. The F1Tenth vehicle is equipped with a front-facing camera that provides visual input for lane following and can achieve scaled speeds comparable to full-scale autonomous vehicles. This platform enables us to evaluate the transferability of our training framework to real-world physical systems. We collected $400$ images in the mini-city and manually annotated the state information for these. We then finetuned the image generator from the CARLA experiments on this dataset to obtain the image generator for the F1Tenth experiments.
@@
For all experiments, we pretrain an anchor controller, which is a dense neural network mapping images to continuous control output. The training is done by imitation learning from a tuned PID controller for the corresponding task, with learning rate of $0.0005$, batch size $256$, and $200$ epochs.
VI-B Image Generator Training and Evaluation
@@
The image generators are implemented as conditional GAN that maps from state information and a dimension $10$ latent vector to greyscale images. The latent vector $z$ is sampled from a uniform distribution $\mathcal{U}(-1,1)$, aiming to capture semantic variations in the driving scene that are not explicitly represented in our state information (e.g., lighting conditions, road textures, environmental elements). The generators output $8\times 16$ greyscale image.
We train the GAN in two stages. First, we train a convolutional GAN. We apply spectral normalization [39] to all discriminator layers and orthogonal regularization [40] to the generator loss function. Both generator and discriminator networks are initialized using orthogonal initialization. They are trained with batch size $128$, learning rate 7e-4, and for $100$ epochs. Then, we distill this GAN into a smaller MLP generator. This simpler network architecture reduces the verification overhead significantly while preserving the state-to-image mapping. During distillation, we minimize a combination of losses $\mathcal{L}_{L1}+\lambda_{adv}\mathcal{L}_{adv}$ where $\mathcal{L}_{L1}$ is the $\ell_{1}$ distance between images generated by the teacher and student generators, $\mathcal{L}_{adv}$ is the adversarial loss from the pre-trained convolutional discriminator.
VI-C Results
Empirical Performance: To evaluate the controllers’ empirical performance, we simulate trajectories starting from $100$ random initial states for $200$ steps. We use the undiscounted cumulative rewards $\sum r_{t}$ as the evaluation metric, where the reward at each step $t$ is defined as $r_{t}=1-\min(1,|d_{t}|/\beta)$, with $d_{t}$ the cross-track error at step $t$ and $\beta$ being a predefined safety threshold.
Figure 1: Empirical performance comparison of controllers.
As shown in Figure 1, the proposed SPVT approach maintains the controller’s empirical performances in all three experiment settings, and it is generally on par with RL controllers in terms of reward maximization.
Safety Guarantee: We collected a dataset of $2000$ i.i.d. initial states. For verification, we used $\alpha,\beta$-CROWN [17, 15]. The safety property we consider is the vehicle not leaving the current lane, equivalently, $|d_{i}|\leq\beta$ for $0\leq i\leq K$. After verifying the controller on this dataset, we can calculate the lower confidence bound for the entire initial state region following Theorem IV.2.
@@ VI-C Results
Figure 2: Semi-Probabilistic Verification (SPV) results: $x$-axis marks the target verification trajectory length ($K$). $y$-axis is the lower bound probability of safety guarantees.
@@
Figure 2 shows the tail bound for safety guarantee over a given range of safety horizon $K$. Compared to all the baseline methods, the proposed SPVT approach significantly increases this probabilistic bound.
VII Conclusion and Limitations

### v2

@@ 4.3 Adaptive Training Data
After we have enough datapoints in $S_{A}$, future training batch $T$ with size $L$ consists of $\lfloor p\cdot L\rfloor$ datapoints from $S_{A}$ and the rest from $S_{0}$, where $p$ is a tunable parameter during training. The portion of datapoints from $S_{0}$ can be sampled uniformly. The portion of datapoints from $S_{A}$ are sampled following the Efraimidis & Spirakis (efraimidis2006weighted) weighted sampling approach to prioritize datapoints that are more difficult (less safe) than others. Specifically, for each datapoint $i$ in $S_{A}$, we first calculate a safety parameter
$\eta^{(i)}$ $=\mathcal{L}_{safety}^{(i)}$ (6)
$+\Big(\max(0,|\overline{\sigma}^{(i)}_{K}|-\beta)+\max(0,|\underline{\sigma}^{(i)}_{K}|-\beta)\Big),$
where the second term further penalizes datapoints whose upper and lower bound is outside of the $\beta,\beta]$ region, where $\beta$ is a predefined safety threshold, and ensure such difficult-to-verify inputs are more likely to be sampled. We then assign a weight $w_{i}=e^{\alpha\cdot\eta^{(i)}}$ (where $\alpha$ is a hyperparameter) to each datapoint $i$ and normalize the weights such that $\sum_{i=1}^{N}w_{i}=1$. Finally we assign each element $x_{i}$ a key $k_{i}=w_{i}^{1/U_{i}}$, where $U_{i}\sim\text{Uniform}(0,1)$. Selecting the top $L$ elements with the highest $k_{i}$ values yields a weighted random sample without replacement, where each element $x_{i}$ is included with probability: $P(x_{i}\in S)=w_{i}/\sum_{j\in\mathcal{B}}w_{j}$.
The full algorithm for data sampling is shown in Algorithm 1. Line $3$ calculates how many datapoints to sample from the buffer based on a tunable parameter $p$; line 484-8 computes the weights and keys following (efraimidis2006weighted); line 9119-11 construct the training batch. This approach enables us to have an adaptive training set which maintains the controller’s average performances while improving its verifiability with respect to datapoints that are more difficult.
@@
5:  Compute $\eta^{(i)}$ according to Equation 6
6:  Compute weight $w_{i}=e^{\alpha\cdot\eta^{(i)}}$
@@
[…]
@@ 4.4 Curriculum Learning
Finally, as in prior work (wu2024verified), we use curriculum learning. Specifically, during training, we first target $K_{i}$-step verified safety. After a series of epochs, we progress to $K_{i+1}$-step verified safety, where $1\leq K_{1}<K_{2}<\cdots<K_{n-1}<K_{n}=K$. This gradual increase in the verification horizon helps improve training stability and enables the controller to learn more complex safety constraints.
5 Experiments
5.1 Experiment Setup
@@
We evaluate our approach in four settings: 1) the X-Plane 11 Flight Simulator, 2) the CARLA Simulator, 3) a mini-city miniature urban physical autonomous driving platform with an F1Tenth racing car, and 4) the Airsim Simulator for drone control. We use two evaluation metrics: 1) empirical performance and 2) lower bound probability for safety guarantee as a function of $K$. We use three baselines for comparison: 1) RESPO (ganai2024iterative), a safe reinforcement learning framework using iterative reachability analysis; 2) SAC-RCBF (emam2022safe), which incorporates safety as a robust-control-barrier-function layer into training; and 3) VSRL (wu2024verified), which guarantees finite-horizon safety by integrating incremental reachability verification into safe reinforcement learning.
@@
Specifically, we first consider the autonomous aircraft taxiing problem using the X-Plane 11 Flight Simulator (xplane11sim). For training the image generator, we collected 20,000 state-image pairs. Each sample consists of an image captured by a forward-facing camera and the corresponding state information (lateral offset $d$ ranging from 10-10 to $10$ meters and heading error $\theta$ between 0.5-0.5 and $0.5$ radians). Data was collected while the aircraft operated at a constant ground speed.
@@
Our second set of experiments consider autonomous lane following using the CARLA Simulator (pmlr-v78-dosovitskiy17a) version 0.9.14. For training the image generator, we sample initial states consisting of the lateral distance $d$ from the lane center, the heading error $\theta$ relative to the lane direction, and the global coordinates $(x,y,z)$ within the CARLA map. The dataset includes trajectories with $d\in[-0.8,0.8]$ meters and $\theta\in[-0.15,0.15]$ radians. Final dataset contains 20,000 state-image pairs collected across different towns (maps) and environmental conditions.
@@
Our third experiment extend the CARLA experiments to a F1Tenth racing car, an open-source 1:10 scale autonomous vehicle, for lane following in a miniature city physical testbed. The F1Tenth vehicle is equipped with a front-facing camera that provides visual input for lane following and can achieve scaled speeds comparable to full-scale autonomous vehicles. This platform enables us to evaluate the transferability of our training framework to real-world physical systems. We collected $400$ images in the mini-city and manually annotated the state information for these. We then finetuned the image generator from the CARLA experiments on this dataset to obtain the image generator for the F1Tenth experiments.
The final problem we consider is drone navigation and obstacle avoidance in Airsim Simulator (airsim2017fsr). The drone is initialized at a fixed hovering position in front of an obstacle observable by its camera. The goal is to learn a smooth control policy while avoid collisions. The image generator is trained on a dataset with $20,000$ samples. The state input consists of the drone’s 3D position $(x,y,z)$ and orientation represented as a quaternion $(q_{w},q_{x},q_{y},q_{z})$. The dataset was collected while the drone was manually controlled by a remote controller.
@@
For all experiments, we pretrain an anchor controller first. For the three path following experiments, the training is done by imitation learning from a tuned PID controller for the corresponding task, with learning rate of $0.0005$, batch size $256$, and $200$ epochs. For the drone experiment, the anchor controller is trained using the standard PPO algorithm with $\gamma$ $0.99$, learning rate $0.0003$ for a total of $200000$ steps.
Then, during safety training, we used supervised training for the path following experiments where we collected a dataset of $30,000$ samples of (state, action) pair. We train the controller for 100 epochs, with batch size $128$ and learing rate $0.0002$. For the drone experiment, we integrated the safety loss directly into RL training using the PPO algorithm. The controller is trained for an additional $200000$ steps with learning rate $0.0002$. Detailed hyperparameters and settings can be found in appendix B.
Figure 1: X-Plane and Carla generator illustrations.
Figure 2: Drone generator illustration.
5.2 Image Generator Training and Evaluation
@@
The image generators are implemented as conditional GAN (cGAN) that maps from state information and a dimension $10$ latent vector to grayscale images; in ablations below we also consider diffusion models for this purpose. The latent vector $z$ is sampled from a uniform distribution $\mathcal{U}(-1,1)$, aiming to capture semantic variations that are not explicitly represented in our state information (e.g., lighting conditions, road textures, environmental elements). The generators output $8\times 16$ images in the three path following experiments, and $64\times 64$ images in the drone experiment.
For the path following experiments, we use a 4-layer fully connected neural network with hidden layers of 256 neurons each, using ReLU activation for the hidden layers and Tanh activation for the output layer. For the drone experiment, we use a more complex convolutional generator; its exact architecture is documented in appendix B. The discriminator in all experiments is a convolutional neural network. During training, we apply spectral normalization (miyato2018spectral) to all discriminator layers and orthogonal regularization (brock2016neural) to the generator loss function. Both generator and discriminator networks are initialized using orthogonal initialization and trained with batch size 128, learning rate 7e-4, for 100 epochs.
Figure 1 and Figure 2 show two ground truth images from the test set for each experimental setting alongside their corresponding generated images. Images from F1Tenth are excluded here as they are visually similar to the Carla images. From visual inspections, the generators perform well for the Carla and Plane experiments, which is largely due to the low-dimensional nature of the images in these settings. For the Drone experiment where we have higher resolution images, on the other hand, while the generated images are less precise, they still capture the essential semantic features needed for effective controller learning.
Figure 3: Empirical validation of Assumption 1.
An important ingredient in our approach is Assumption 1, that for true image inputs we can find a latent representation $z$ as input to the generator that results in its close approximation. We validate this empirically for $8$ choices of latent dimensions: $2,4,6,8,10,16,32,64$. We evaluate the generators on a separate test set: for each ground-truth image, we perform gradient-based optimization to find the best latent variable $z$ that minimizes the mean squared error (MSE). Notably, since this is not guaranteed to find a globally optimal $z$, our results provide a pessimistic evaluation (i.e., it is possible that such a $z$ exists even if we fail to find it using gradient-based optimization).
As shown in Figure 3, as latent dimension increases, the averaged MSE rapidly decreases, i.e., the generator $g$ better approximates the unknown $h$ in all four experiment settings we consider in this work. Moreover, we can find effective approximations of the true image inputs for even relatively small latent dimension.
Figure 4: Empirical performance comparison of controllers. Box plots show the distribution of episode rewards over 100 evaluation episodes.
@@
Figure 5: Semi-Probabilistic Verification (SPV) results: $x$-axis marks the target verification trajectory length ($K$). $y$-axis is the lower bound safety probability.
5.3 Results
Empirical Performance
To evaluate the controllers’ empirical performance, we simulate trajectories starting from $100$ random initial states for $200$ steps. We use the undiscounted cumulative rewards $\sum r_{t}$ as the evaluation metric. For the path following tasks (X-Plane, Carla, and Physical Minicity), the reward at each step $t$ is defined as $r_{t}=1-\min(1,|d_{t}|/\beta)$, where $d_{t}$ is the cross-track error at step $t$ and $\beta$ is a predefined safety threshold. We set $\beta=1.0$m for Carla and Physical Minicity evaluations, and $\beta=12.5$m for X-Plane evaluations.
For drone control in Airsim, we used the same reward function in both training and evaluation. The reward includes components for forward velocity, attitude stability, and control smoothness. The complete reward function is provided in appendix A.
As shown in Figure 4, the proposed SPVT approach maintains strong empirical performance across all experimental settings. Notably, in the Drone experiment, which has the highest-dimensional image input, SPVT demonstrates significantly superior performance compared to all safe training baselines. This result suggests that SPVT is the only scalable safe training method capable of handling high-dimensional image inputs. It is worth noting that VSRL training fails entirely in the Drone experiment because VSRL requires formal verification over the controller’s complete input space, which becomes computationally infeasible for high-dimensional image data.
Safety Guarantee
For the path following tasks, the safety property we consider is the vehicle not leaving the current lane, equivalently, $|d_{i}|\leq\beta$ for $0\leq i\leq K$. For the drone experiment, safety properties include altitude bounds ($z\in[0.2,15]$ meters) and attitude limits (roll and pitch angles $\leq 1.2$ radians).
To obtain the lower-bound safety probability under our semi-probabilistic verification (SPV) framework, we collected a dataset of $2000$ initial states sampled i.i.d. from the initial state distribution. We used $\alpha,\beta$-CROWN (zhang2018efficient; xu2020automatic; xu2021fast) to verify the model iteratively. At each step, $\alpha,\beta$-CROWN computes bounds on the model output, which, after dynamics calculations, gives the reachable state regions at the next timestep. We then check whether this state region intersects with unsafe states. We iterate this forward reachability analysis for $K$ steps, obtaining empirical verified safety result, i.e. the proportion of initial states in our dataset that are verified safe at each step. These empirical results are then used in Theorem 2 to compute formal lower-bound safety probabilities over the entire initial state space. We set the confidence parameter $\delta=0.05$ for all experiments.
@@ 5.3 Results
Figure 5 shows the resulting safety probability lower bounds as a function of horizon $K$. Compared to all baseline methods, the proposed SPVT approach significantly increases these probabilistic safety bounds.
Ablations
Figure 6: Ablation study comparing cGAN vs. diffusion-based state-to-image mapping in X-Plane. Left: controller performance over 100 episodes. Right: SPV verified safety probability bounds.
A natural alternative to our cGAN-based state-to-image mapping is to use diffusion-based image generation. In this case, $g(s,z)=z-\phi(x,z)$, where the diffusion model $\phi(x,z)$ predicts the noise in the randomly generated input. To provide finer control over the dimension of $z$, we make use of an autoencoder $\psi(z)$ that takes randomly generated vector $z$ of a given dimension as an input, and outputs a decoded noisy image $z^{\prime}=\psi(z)$ with the dimension of the image, so that $g(s,z)=\psi(z)-\phi(x,\psi(z))$. To evaluate this design choice, we implement a diffusion-based variant of SPVT and compare it against our cGAN approach on both controller performance and safety verification. Figure 6 shows results in the X-Plane environment. We find that controller trainied in the cGAN setup achieve higher rewards and maintain better safety probability bounds. Ablation results in other environments, along with more detailed analysis of the observed difference, are provided in Appendix C.
6 Related Work
Formally verifying the safety of vision-based controllers is extremely challenging. Traditional verification tools (katz2017reluplex; huang2017safety; ehlers2017formal; sinha2017certifying) are too computationally demanding for this purpose. Recent progress has enabled verification of larger neural networks through over-approximation (zhang2018efficient; gowal2018effectiveness; xu2021fast), and abstract interpretation (gehr2018ai2; singh2019abstract; katz2019marabou). However, verifying vision-based controllers remains challenging due to high dimensionality, as well as the difficulty of defining safety properties in the image space.
A practical approach to verify vision-based controllers is by approximating the perception module. This reduces verification complexity by projecting the problem from image space to lower-dimensional state space (katz2022verification; cai2024scalable; hsieh2022verifying). These methods, however, only focus on post-hoc verification and cannot be integrated into the training process for learning a safer controller. For a more comprehensive survey on the verification of vision-based controllers, see (vision_based_control_verification_survey_2024).
Safe reinforcement learning encodes safety based on the constrained Markov decision processes (altman2021constrained) but it is often too soft to enforce strict safety constraints (wang2023). On the other hand, control-theoretic methods such as barrier functions (dawson2022safe; emam2022safe; barriernet_2023), and reachability analysis (yu2022reachability; ganai2024iterative) are more powerful but they have limited scalability to high-dimensional system such as vision-based controllers. Recently, some works extend these methods to the image space (V-CBF_Abdi_ICRA2023; NeRF_based_CBF_ICRA_2023; yuxuan_2024_icra; Ihab_2024_L4DC). However, they function more like a safety check around the controller, giving yes/no answers for passed in reference control, without providing formal safety guarantees about the controller itself.
7 Conclusion and Limitations